OpenStudy (anonymous):

determine dy/dx by implicit differentiation 5x^2y-sin(x^2+y^2)-2x=0

sam (.sam.):

Where are you stucked?

OpenStudy (anonymous):

by implicit differentiation... kind of wondering if this is the chain rule...

OpenStudy (anonymous):

i'm not too good with this stuff

OpenStudy (mathmale):

Apply the derivative operator to the implicit equation that you've typed in:\[\frac{ d }{ dx }(5x ^{2}y-\sin(x^2+y^2)-2x=0)\]

OpenStudy (mathmale):

Implicit diff'n almost always involves the chain rule, but is not the chain rule. the first term inside the parentheses is 5x^2*y, which we should recognize as being a product. so we'll have to apply the product rule here, and, by the way, also the chain rule on that y. Can you find \[\frac{ d }{ dx }5x^2y?\]

OpenStudy (anonymous):

i think that is 10xy

OpenStudy (mathmale):

If you had applied the product rule, you would have obtained a result with two terms. \[\frac{ d }{ dx }5x^2y=5(x^2\frac{ dy }{ dx }+y(2x))\]Please look this over carefully and determine how much sense (or how little) it makes to you.

OpenStudy (mathmale):

Some people write the product rule as (uv)' = uv' + vu' which is nice and short.

OpenStudy (anonymous):

okay that makes sense above. now what about the -2x? does that end up becoming, 5(2xy+x^2(dy/dx)-2?